OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left A number line with a closed circle on 6 and shading to the right

OpenStudy (anonymous):

help

OpenStudy (anonymous):

sooo its \[3(2x-1)\ge 4(2x-3) -3\] \[6x-3 \ge 8x-12-3\] \[6x \ge 8x -12\] \[-2x \le -12\] \[x \ge 6\] x less than or greater to 6 A number line with a closed circle on 6 and shading to the left

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

can you help me in other one

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Uhm o.o okay

OpenStudy (anonymous):

Solve for x. -ax + 3b > 5 x > the quantity 3 times b minus 5 all over negative a x > the quantity 5 minus 3 times b all over negative a x < the quantity 3 times b plus 5 all over negative a x < the quantity negative 3 times b plus 5 all over negative a

OpenStudy (anonymous):

x < the quantity 3 times b plus 5 all over negative a